OpenStudy (anonymous):

2x-4y=8 state slope and y-intercept

OpenStudy (lgbasallote):

do you know how to solve for y in your equation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

isolate the variable.. thus get y alone on once side and the x term and the constant onthe other side

OpenStudy (anonymous):

-4y = 8 - 2x

OpenStudy (anonymous):

get rid of the -4 using division. can you solve from here?

OpenStudy (anonymous):

Add 4 to both sides . the two 4's cross out

OpenStudy (anonymous):

-4y/-4=-2x+8/-4 y=1/2x-2 is that right?

OpenStudy (anonymous):

y=mx+b where m is the slope and b is the y intercept

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

yay thanks :D

OpenStudy (anonymous):

good luck